There is 15 liters of a 12 percent dye solution. How much water must be evaporated from this mixture to obtain a solution that is 20 percent dye? Assume that the total number of dye is not affected by the process of evaporation.
Answer by Theo(13342)

.12 * 15 = 1.8 liters of dye.
let x = total amount of the new solution.
then 1.8 / x = .2
solve for x to get x = 1.8 / .2 = 9
the new solution contains 9 liters of both dye and water.
the water in the new solution is equal to 9 - 1.8 = 7.2 liters.
the water in the old solution is equal to .88 * 15 = 13.2 liters.
the amount of water to be removed from the old solution is equal to 13.2 - 7.2 = 6 liters.
your original solutions contained:
1.8 liters of dye.
13.2 liters of water.
total liters of original solution is equal to 15.
your new solution contains:
1.8 liters of dye.
7.2 liters of water.
total liters of new solution is equal to 9.